two minor league baseball players got a total of 308 hits. Washington has 8 more hits than Sanchez. find the number of hits for each player.
1
Expert's answer
2013-07-22T08:32:16-0400
Let Sanchez has x hits. Then Washington has x+8.
We can construct the following equation:
x+x+8=308
2x=300
x=150 that's how many hits Sanchez has.
Washington has x+8=150+8=158.